OpenStudy (anonymous):

If the radius of A is 5 units, calculate the area

OpenStudy (sleepyhead314):

Hello @Korean and Welcome to OpenStudy! :) area of a circle = pi * (radius)^2 so area = pi * (5) * (5) = ?

OpenStudy (anonymous):

If the radius of A is 5 units, calculate the area.

OpenStudy (sleepyhead314):

you mean the area of the shaded region?

OpenStudy (anonymous):

Umm I think so

OpenStudy (sleepyhead314):

well that would be area of the circle minus the area of the triangles

OpenStudy (anonymous):

Umm but I don't know how to measure the area of the triangles